OpenStudy (anonymous):

when 8 consecutive integers are multiplied, their product is 0. What would be their largest possible sum?

OpenStudy (anonymous):

28

myininaya (myininaya):

we have x*(x+1)*(x+2)*(x+3)*(x+4)*(x+5)*(x+6)*(x+7)=0 these are 8 consecutive numbers whose product is 0 that means one of these have to be 0 in order for the product to be 0 and since we want the largest possible sum we say x=0 so x+1=0+1=1 and x+2=0+2=2 ... x+7=0+7=7 so we have 0+1+2+3+4+5+6+7=? you fill in this blank

myininaya (myininaya):

yes and 28 is right gj!